1,00 liter of solution A was prepared by dissolved 15,40 gram of 3,4−DHBA acid and 15,40 gram 3,5−DHBA acid on distilled water.

1,0 liter of solution B was prepared by dissolved 30,80 gram mixture of 3,4−DHBA acid and 3,5 – DHBA acid. pH of solution B is 2.40.

a) Calculate pH value of solution A.

b) How many gram of each acid (3,4−DHBA acid and 3,5 – DHBA acid) was used to prepare 1,0 liter of solution B.

c) Describe the method to prepare 1,00 liter of buffer solution with pH = 4,11 by using distilled water, solid 3,4−dihydroxybenzoic acid and solid sodium 3,4−dihydroxybenzoate. In this buffer solution: Ca + Cb = 1,00 M (Ca is 3,4−dihydroxybenzoic acid concentration; Cb is sodium 3,4−dihydroxybenzoate concentration).

Given by: M(DHBA) = 154 gam.mol−1 .

pKa 3,4− dihydroxybenzoic acid (3,4−DHBA) 4,48

pKa 3,5− dihydroxybenzoic acid (3,5−DHBA) 4,0


1
Expert's answer
2021-08-11T05:21:46-0400

Molar mass of 3,4−DHBA = 184.15

= 1540/184.15

= 8.363 moles

Molarity = 8.363/100

= 0.0836

PH = 14 - log(0.0836

= 14+0.778

= 14.778